the width and length of a rectangle are in the ratio of 5/4.How wide in the rectangle if the length is 34 cm more the width?

Answer by geetha_rama(94)

Let L be the length of the rectangle and W be the width
L/W = 5/4
=>4L = 5W
Length is 34 cm more than width
L = W+ 34
=> 4(W+34) = 5W
136 = W
Width = 136 cm
Length = W+ 34 = 170 cm
Check : L/W = 170/134 = 17*10/17*8 = 5/4